car is traveling at constant speed of 20.8 m/s when the car sees sign that is 50.0 meters infront dicides to break the pedal for a constant acceleration of 4.80 m/s2 to reach in a rest mode. What is the cars reaction time?
1
Expert's answer
2019-03-04T10:46:45-0500

The distance after he decides to break the pedal


"d=v^2\/(2a)=(20.8)^2\/(2(4.8))=45.07 m."

The reaction time is


"t=(50.0-45.07)\/20.8=0.237 s."